The dimensions of a retangle are 5cm by 13m. what are the dimensions of a similar rectangle with exactly double the area?
Answer by jim_thompson5910(35256)

A = LW

A = 5*13

A = 65


Double the area to get 65*2 = 130


A = LW

130 = xy

y = 130/x

--------------

x/y = 13/5


y = 13x/5

-------------

y = 130/x

13x/5 = 130/x

13x^2 = 130*5

13x^2 = 650

x^2 = 650/13

x^2 = 50

x = sqrt(50)

x = 5*sqrt(2)

x = 7.0710678
---------------------
y = 130/x

y = 130/7.0710678

y = 18.384776



So the dimensions are roughly 7.0710678 by 18.384776